OpenStudy (anonymous):

Will give medal and become fan!! How do u convert the repeating, nonterminating decimal to a fraction? Problem is 0.1515...

OpenStudy (mathmath333):

\(\large 0.151515....=\dfrac{15}{99}\)

OpenStudy (anonymous):

thanks so much

OpenStudy (mathmath333):

note that i have placed 9 two times because 15 is a two digit number

OpenStudy (mathmath333):

and u can further simplify \(\large 0.151515....=\dfrac{15}{99}=\dfrac{5}{33}\)